# Dynamical complexity of non-Gaussian many-body systems with dissipation

### v1

@@ Abstract
Finally, we consider if a version of Theorem 2 holds for fermions as well i.e., if noise rates larger than the Gaussian inter-site couplings result in separability at all times, as might be intuitively expected. However, we answer this question in the negative: in the Supplemental Material we show that for fermionic systems, in stark contrast with bosons, no matter how high the rate of particle loss $(\kappa\xi^{(1)})$ and incoherent particle gain $(\kappa\xi^{(2)})$ is, the system does not remain separable at all times and can exhibit entanglement at short times. In fact, we show that this short-time non-separability, irrespective of the decay rate $\kappa$, holds not only if we consider separability with respect to all observables [61], but also if we consider a weaker notion of separability with respect to only parity-conserving observables [62]. However, this result does not rule out separability at longer times, or exclude an efficient classical simulation of the fermionic model that does not rely on its separability.
Conclusion. We have characterized the classical complexity of simulating the continuous-time evolution of fermionic and bosonic systems as a function of the noise, Gaussian and non-Gaussian interaction strengths. Furthermore, in the Supplemental material we present a slightly more general model, where non-Gaussian inter-site interaction terms are also allowed [50]. Future directions include the study of other experimentally relevant models, such as non-Markovian quantum systems. We leave for future works the question of whether fermionic systems without dephasing noise can encode universal computations with noise rates much larger than the non-Gaussianity.

### v2

@@ Abstract
Finally, we consider if a version of Theorem 2 holds for fermions i.e., do noise rates larger than the Gaussian inter-site couplings result in separability at all times. We answer this question in the negative: in the Supplement we show by analyzing few-mode fermionic models that, in contrast with bosons, no matter how high $\kappa_{2},\kappa_{3}$ are, the system does not remain separable at all times and can exhibit entanglement at time-scales $\sim\min(\kappa_{2}^{-1},\kappa_{3}^{-1})$. In fact, this short-time non-separability holds not only if we consider separability with respect to all observables [61], but also if we consider a weaker notion of separability with respect to only parity-conserving observables [62]. However, this result does not rule out separability at longer times or other routes to classical simulation.
Conclusion and outlook. We have characterized the classical complexity of simulating the continuous-time evolution of fermionic and bosonic systems as a function of the noise, Gaussian and non-Gaussian interaction strengths. Future theoretical directions include the study extending our results to non-Markovian models of dissipation.
The models considered in this paper can be experimentally realised in several platforms. The bosonic model can be implemented in superconducting systems where the Gaussian Hamiltonian can be controlled by designing capacitive couplings between different qubits and the single-site non-Gaussianity by the nonlinear Josephson potential in the qubit [28, 29]. We can also use cold bosonic atoms in optical lattices where the strength of both the Gaussian and the non-Gaussian Hamiltonians can be controlled by tuning the optical lattice potential [26, 27]. The fermionic model can be implemented either with cold atoms in optical lattices by using a fermionic species of atoms [23, 24, 25], or in solid-state systems such Moiré superlattices hosting trions [63, 64, 65]. Since all of these systems will have intrinsic particle loss, gain, and dephasing, tuning the parameters in the Gaussian and non-Gaussian Hamiltonians could allow us to access the parameter regimes in Theorems 1 and 2. A major challenge in experimentally verifying the threshold behavior predicted by Theorems 1 and 2 would be verifying the presence (or absence) of entanglement/non-Gaussianity in $\rho(t)$. While this could be hard to do for large systems, we remark that the difference in the threshold behavior in fermionic and bosonic models that we described can be understood even with systems with few ($\leq 4$) fermionic or bosonic modes, which is well within the regime where a full state tomography can already be performed.